Two angles are supplementary . the measure of one angle is 10 more than three timed the other . what is the measure of the obtuse angle

Answer by Aldorozos(172)

x+3x+10 = 180
4x =170 x=170/4 = 42.5 If one angle is 42.5 the other angle is 180-42.5 = 137.5
137.5 which is greater than 90 is the obtuse angle.